OpenStudy (anonymous):

a motorist travelling with a constant speed of 15 m/s passes a school zone sign where the speed limit is 10 m/s. Just as the motorist passes, a police officer on a motorcycle starts off at the same school zone sign in pursuit with constant acceleration of 20 m/s^2.

OpenStudy (anonymous):

Well I assume you are asking when he catches the motorist Lets start off by taking the board sign as the origin Now lets say they meet after time t Dist.travelled by motorist in time t=10t By policeman=\[20(t^2)\] Now equate them because that is where they meet You get two answers t=0 and t=2

OpenStudy (anonymous):

ah no thats wrong

OpenStudy (anonymous):

Go on explain why

OpenStudy (anonymous):

t=0.75s v=v0+at

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

velocity = initial velocity + acceleration*time

OpenStudy (anonymous):

So what the motorist didn't have any acceleration.And I equated the distance travelled by both now the velocities

OpenStudy (anonymous):

his is constant

OpenStudy (anonymous):

Yeah so what's you point

OpenStudy (anonymous):

lol oops forgot to post this part LOL (a)How much time elapses before the officer catches up with the motorcycle? (b)What is the total distance that each have travelled at that point?

OpenStudy (anonymous):

Well he catches after 2 seconds at a distance 20 m later

OpenStudy (anonymous):

just checking answer...i forgot to post a bit of the question (a)How much time elapses before the officer catches up with the motorcycle? (b)What is the total distance that each have travelled at that point?

OpenStudy (anonymous):

5.65m?

OpenStudy (anonymous):

sorry 5.625

OpenStudy (anonymous):

and 0.75s

OpenStudy (anonymous):

after t seconds the motorist goes 15t meters, i believe the formula for distance is d=v_ot+1/2at^2

OpenStudy (anonymous):

But the policeman's initial velocity was 0

OpenStudy (anonymous):

Oh I forgot and kept initial velocity for biker as 10 instead of 15

OpenStudy (anonymous):

Anyway just change the numbers

OpenStudy (anonymous):

yes, so 15t=1/2*20t^2

OpenStudy (anonymous):

Yeah that'd give you the answer

OpenStudy (anonymous):

but t=.75s? and so answer is 5.625?

OpenStudy (anonymous):

t=1.5s

OpenStudy (anonymous):

total distance travelled is 15(1.5)=22.5m

OpenStudy (anonymous):

ohhh so you use a quadratic to find t in d=v_ot+1/2at^2

OpenStudy (anonymous):

i was using v=vo+at

OpenStudy (anonymous):

so if there is distance involved you have to use that formula...

OpenStudy (anonymous):

first you have to equate 15t with v_ot+1/2at^2, solve for t, then solve for d

OpenStudy (anonymous):

distance traveled by motorist in t seconds is 15t, distance travelled by cop in t seconds is 10t^2,

OpenStudy (anonymous):

haha thanks